Alice chicken coop(cage) has a total volume of 624 cubic feet. Use the fact that 1 foot is approximately equal to 0.3048m to convert this volume to cubic meter.
1
Expert's answer
2020-10-28T16:53:19-0400
"1\ ft=0.3048\ m"

"1\ ft^3 =(0.3048)^3 \ m^3"

"V=624\ ft^3=624(0.3048)^3 m^3="

"=17.669712273408 \ m^3"

"V\approx17.669712\ m^3"